OpenStudy (anonymous):

Find the distance between (4 sqrt 2, -3) and (5 sqrt 2, 2). express answer in simplest form. its a multiple choice question A) sqrt 163 B) 81 sqrt 2 C) 3 sqrt 3 D) sqrt 3

OpenStudy (anonymous):

considering dat u know the distance formula, the answer is:C and d'u know it?

OpenStudy (anonymous):

confused....

OpenStudy (anonymous):

its C. got it

OpenStudy (anonymous):

okie, now the 2 points are \[(4\sqrt{2},-3) and (5\sqrt{2},2)\] distance formula states: distance between 2 points (x1,y1) and (x2,y2) is : sqrt((x2-x1)^2+(y2-y1)^2); so we do the same: sqrt((5root2-4root2)^2+(2-(-3))^2) now 5root2-4root2=root2 this is normal subtraction, take root2 out u get, root2(5-4) so squaring it u hav 2+5^2=27 which is 3^2*3 and so C.

OpenStudy (anonymous):

thanks